# Accelerated Parallel Tempering via Neural Transports

### v3

@@ 4 Analysis of Accelerated PT
We analyse how the performance, measured by the round trip observed after $T$ iterations, $R_{T}$, scales with increasing parallel chains $N$ and acceleration time $K$. It is equivalent to analyse the round trip rate $\tau:=\lim_{T\to\infty}\mathbb{E}[R_{T}]/T$, defined as the expected percentage of PT iterations, where a round trip occurs (Katzgraber et al., 2006; Lingenheil et al., 2009), which coincides with slope in Figure˜1 (Right). To derive theoretical insights independent of the problem-specific local exploration move, we make an efficient local exploration assumption analogous to Syed et al. (2021; 2022); Surjanovic et al. (2024), which assumes the work done by forward and backward accelerated paths is independent across chains and PT iterations.
@@
Proposition˜2 relates the round trip rate to the rejections, extending Syed et al. (2021, Corollary 2).
@@
We now fix $N$ and use Proposition˜2 to study how the round trip rate scales with $K$. We focus on a special case where ${\mathbb{P}}_{K}^{n-1}$ and ${\mathbb{Q}}_{K}^{n}$ arise as $K$-step discretisations of an underlying stochastic differential equation (SDE) which bridge $\pi^{n-1}$ and $\pi^{n}$ with path measure ${\mathbb{P}}^{n-1}_{\infty}$ and ${\mathbb{Q}}^{n}_{\infty}$ as $K\to\infty$.
@@
Under appropriate conditions on the drifts of the SDE (Section˜B.2), as $K\to\infty$, $\tau({\mathbb{P}}^{0:N-1}_{K},{\mathbb{Q}}^{1:N}_{K})$ converges to $\tau({\mathbb{P}}^{0:N-1}_{\infty},{\mathbb{Q}}^{1:N}_{\infty})$ and $r({\mathbb{P}}_{K}^{n-1},{\mathbb{Q}}_{K}^{n})\leq r({\mathbb{P}}^{n-1}_{\infty},{\mathbb{Q}}^{n}_{\infty})+\mathcal{O}(\frac{1}{\sqrt{K}})$.
The main appeal of accelerated PT comes from the fact that for good neural network approximations
@@
Thus, increasing $K$ can dramatically reduce the rejection rate and increase the round trip rate in challenging scenarios. In practice, we find that for moderate values of $K$, this benefit overweights the increased computational cost as seen in Table˜1.
@@
To understand how accelerated PT scales with increased parallelism, as $N\to\infty$ for a fixed number of acceleration steps $K$. For each $N$, we assume $\pi^{n}$ discretizes a continuous annealing path $\pi^{\beta}$ interpolating between reference and target at $\beta=\beta_{n}$ for some annealing schedule $0=\beta_{0}<\dots<\beta_{n}=1$. Additionally we assume ${\mathbb{P}}^{n-1}_{K}$ and ${\mathbb{Q}}^{n}_{K}$ discretize the mutually absolutely continuous forward/backward accelerators ${\mathbb{P}}^{\beta,\beta^{\prime}}_{K}$ and ${\mathbb{Q}}^{\beta,\beta^{\prime}}_{K}$ between $\pi_{\beta}$ and $\pi_{\beta^{\prime}}$ at $\beta=\beta_{n-1}$ and $\beta^{\prime}=\beta_{n}$ respectively with work function $W^{\beta,\beta^{\prime}}_{K}$.
@@
Suppose ${\mathbb{P}}^{\beta,\beta^{\prime}}_{K}$ and ${\mathbb{Q}}^{\beta,\beta^{\prime}}_{K}$ are sufficiently regular and satisfy Assumptions 2–4 in Section˜B.3. As $N\to\infty$ if $\max_{n\leq N}|\beta_{n}-\beta_{n-1}|=O(N^{-1})$, then $\sum_{n=1}^{N}r({\mathbb{P}}^{n-1}_{K},{\mathbb{Q}}^{n}_{K})$ converges to $\Lambda_{K}$ and $\tau({\mathbb{P}}^{0:N-1}_{K},{\mathbb{Q}}^{1:N}_{K})$ converges to $\bar{\tau}_{K}=(2+2\Lambda_{K})^{-1}$, where $\Lambda_{K}$ equals,
@@
and $\dot{W}^{\beta}_{K}:\mathcal{X}^{K+1}\to\mathbb{R}$ is the partial derivative with respect to $\beta^{\prime}$ of $W^{\beta,\beta^{\prime}}_{K}$ at $\beta^{\prime}=\beta$.
Theorem˜2 shows that as $N\to\infty$, the round trip rate is controlled by the global barrier for APT, $\Lambda_{K}$, and can be approximated by rejection rates obtained in Algorithm 1. Notably, APT observes a sharp deterioration in round trip when $N\ll\Lambda_{K}$, begins to stabilise when $N\approx\Lambda_{K}$ and observes a marginal improvement when $N\gg\Lambda_{K}$. As discussed in Proposition˜3, for a reasonable choice of accelerators we expect $\Lambda_{K}$ to decrease with $K$, which suggests APT can substantially improve compared to classical PT ($K=0$) for challenging problems where $N\approx\Lambda_{0}>\Lambda_{K}$. Finally, we explore the trade-off between increasing parallelism and increasing acceleration time in Appendix˜B.
5 Design Space for Diffusion PT
@@ 5.1 Normalising Flow Accelerated PT
Let $T^{n}:\mathcal{X}\to\mathcal{X}$ be such a diffeomorphism. We denote NF-APT as APT with forward and backward accelerators to deterministically transport by normalising flow $T^{n}$ and its inverse, respectively, motivated by the use of normalising flows to map between annealing distributions in Arbel et al. (2021) and Matthews et al. (2022) in the context of SMC samplers (Del Moral et al., 2006),
@@
Then the work functional equals,
$W^{n}_{1}(x_{0},x_{1})=U^{n}(x_{1})-U^{n-1}(x_{0})-\log|\det\nabla T^{n}(x_{0})|,\quad x_{1}=T^{n}(x_{0}).$
We can parametrise the normalising flow $T^{n}$ by a neural network trained to optimise the symmetric KL-divergence from Equation˜4, $\mathcal{L}(T)=\sum_{n=1}^{N}\mathrm{SKL}({\mathbb{P}}^{n-1}_{K},{\mathbb{Q}}^{n}_{K})$. See Section˜C.1 for details.
@@ 5.2 Controlled Monte Carlo Diffusions Accelerated PT
Another common transport for sampling is based on the escorted Jarzynski equality (Vaikuntanathan & Jarzynski, 2008). One way to realise this in practice with neural networks is via Controlled Monte Carlo Diffusions (CMCD) (Vargas et al., 2024). Suppose $\mathcal{X}\subseteq\mathbb{R}^{d}$ and for $s\in[0,1]$ we have $b^{n}_{s}\in\mathbb{R}^{d}$, $\sigma^{n}_{s}\in\mathbb{R}_{+}$, and $U^{n}_{s}=(1-\phi_{s}^{n})U^{n-1}+\phi_{s}^{n}U^{n}$, where $\phi_{s}^{n}\in[0,1]$ is a monotonically increasing in $s$ and $\phi^{n}_{0}=0,\phi^{n}_{1}=1$. For a fixed $K$, let $s_{k}=k/K$ and $\Delta s_{k}=1/K$ be the uniform discretisation of the unit interval. We denote CMCD-APT as APT with $k$-th forward transition kernel,
$P^{n-1}_{k}(x_{k-1},\mathrm{d}x_{k})=\mathcal{N}(x_{k-1}-(\sigma^{n}_{s_{k-1}})^{2}\nabla U^{n}_{s_{k-1}}(x_{k-1})\Delta s_{k}+b^{n}_{s_{k-1}}(x_{k-1})\Delta s_{k},2(\sigma^{n}_{s_{k-1}})^{2}\Delta s_{k}),$
and the backward transition,
$Q^{n}_{k-1}(x_{k},\mathrm{d}x_{k-1})=\mathcal{N}(x_{k}+(\sigma_{s_{k}}^{n})^{2}\nabla U^{n}_{s_{k}}(x_{k})\Delta s_{k}+b^{n}_{s_{k}}(x_{k})\Delta s_{k},2(\sigma_{s_{k}}^{n})^{2}\Delta s_{k}).$
The corresponding work function for CMCD-APT equals,
$W^{n}_{K}(x_{0:K})=U^{n}(x_{K})-U^{n-1}(x_{0})+\sum_{k=1}^{K}\log P^{n-1}_{k}(x_{k-1},x_{k})-\sum_{k=1}^{K}\log Q^{n}_{k-1}(x_{k},x_{k-1}),$ (5)
where $P^{n-1}_{k}(x_{k-1},x_{k})$ and $Q^{n}_{k}(x_{k},x_{k-1})$ are the Gaussian densities of the forward and backwards kernel transitions with respect to the Lebesgue measure. We learn the neural transport $b^{n}_{s}$, the time schedule $\phi^{n}_{s}$ and the diffusion coefficient $\sigma^{n}_{s}$. We use a symmetrised KL-divergence loss $\mathcal{L}(b_{s},\phi_{s},\sigma_{s})=\sum_{n=1}^{N}\mathrm{SKL}({\mathbb{P}}_{K}^{n-1},{\mathbb{Q}}^{n}_{K})$, where we also learn $\phi_{s}^{n},\sigma^{n}_{s}$ following Geffner & Domke (2023) [^2] We note that in standard CMCD, we cannot use symmetrised KL-divergence as the objective, as we only have access to data from one side.. While in theory, once the vectorised field $b_{s}^{n}$ is perfectly learned, any $\sigma^{n}_{s}$ can be used, we found that learning $\sigma^{n}_{s}$ can significantly stabilise the training and enhance the results. Note that other divergences and losses can be used, such as (Máté & Fleuret, 2023; Richter & Berner, 2023; Albergo & Vanden-Eijnden, 2024). We discuss the form of the transition kernels, the work, and loss in the limit at $K\to\infty$ in Section˜C.2.
@@ 5.3 Diffusion Accelerated PT
Given $\mathcal{X}=\mathbb{R}^{d}$, the Variance-Preserving (VP) diffusion model (Ho et al., 2020; Song et al., 2020) is defined by the SDE $dY_{s}=-\gamma_{s}Y_{s}\mathrm{d}s+\sqrt{2\gamma_{s}}\mathrm{d}{W}_{s}$, with $s\in[0,1]$, $Y_{0}\sim\pi$, and $\gamma_{s}$ is a rate function from $[0,1]$ to $\mathbb{R}^{+}$. The time-reversal SDE $(X_{s})_{s\in[0,1]}=(Y_{1-s})_{s\in[0,1]}$ has the form $dX_{s}=[\gamma_{1-s}X_{s}+2\gamma_{1-s}\nabla\log\pi_{s}^{\text{VP}}(X_{s})]\mathrm{d}s+\sqrt{2\gamma_{1-s}}\mathrm{d}{W}_{s}$ where $\pi_{s}^{\text{VP}}$ is the density of $Y_{1-s}$ and $X_{0}\sim\pi_{0}^{\text{VP}}={\mathcal{N}}(0,\mathrm{I})$ is close to a standard Gaussian in the common scenario that $\int_{0}^{1}\gamma_{s}\mathrm{d}s\gg 1$. The score $\nabla\log\pi_{s}^{\text{VP}}$ is unknown but can be learned by score-matching objectives.
We introduce for $s\in[0,1]$ an energy-based model (Salimans & Ho, 2021) with potential $\tilde{U}^{\text{VP}}_{s}:\mathcal{X}\mapsto\mathbb{R}$ and corresponding annealing distributions $\tilde{\pi}^{\text{VP}}_{s}\propto\exp(-\tilde{U}^{\text{VP}}_{s})$ to approximate $\pi_{s}^{\text{VP}}$ while satisfying the boundary conditions $\tilde{\pi}^{\text{VP}}_{0}={\mathcal{N}}(0,\mathrm{I})$ and $\tilde{\pi}^{\text{VP}}_{1}=\pi$. For some schedule $0=s_{0}<\cdots<s_{N}=1$, we consider the annealing path $\pi^{n}=\tilde{\pi}^{\text{VP}}_{s_{n}}$. For a fixed $K$, we denote Diff-APT as APT with forward and backward kernels respectively given by some $K$-step integrator of the time-reverse SDE and the forward SDE between $s_{n-1}$ and $s_{n}$. For instance, the Euler integrator with step-size $\delta_{n}=(s_{n}-s_{n-1})/K$ and interpolating times $s_{n,k}=s_{n-1}+k\delta_{n}$ yields:
$P_{k}^{n-1}(x_{k-1},\mathrm{d}x_{k})$ $=\mathcal{N}(x_{k-1}+\delta_{n}\gamma_{1-s_{n,k}}x_{k-1}-2\gamma_{1-s_{n,k}}\nabla\tilde{U}_{s_{n,k},\theta}^{\text{VP}}(x_{k-1}),2\delta_{n})$ (6)
$Q_{k-1}^{n}(x_{k},\mathrm{d}x_{k-1})$ $=\mathcal{N}(x_{k}-\delta_{n}\gamma_{1-s_{n,k}}x_{k},2\delta_{n}).$ (7)
The resulting work function $W^{n}_{K}$ coincides with Equation˜5 from CMCD-APT, replacing the forward backward transitions $P^{n-1}_{k}$ and $Q^{n}_{n-1}$ for CMCD-APT with Equation˜6 and Equation˜7. Further, we parametrise $\tilde{U}_{s}^{\text{VP}}$ by a neural network trained to minimise the standard score matching objective. Since we cannot access samples from $X_{1}\sim\pi$, we iteratively train on approximate samples from $\pi$ obtained by sampling from Diff-APT. We stress that a parallelised implementation of Diff-APT allows for generating a sample from $\pi$ with a similar effective cost as a single discretisation step solving the time-reversal SDE. For further details, see Section˜C.3.

### v4

@@ 4 Analysis of Accelerated PT
We analyse how the performance, measured by the round trips $R_{T}$ observed after $T$ iterations, scales with increasing parallel chains $N$ and acceleration time $K$. This is equivalent to analysing the round trip rate $\tau:=\lim_{T\to\infty}\mathbb{E}[R_{T}]/T$ defined as the expected percentage of PT iterations where a round trip occurs (Katzgraber et al., 2006; Lingenheil et al., 2009). We use the slope in Figure˜1 (Right) to illustrate this quantity. To derive theoretical insights independent of the problem-specific local exploration moves, we employ an efficient local exploration assumption analogous to Syed et al. (2021; 2022); Surjanovic et al. (2024), which assumes the weight of forward and backward accelerated paths is independent across chains and PT iterations.
@@
Proposition˜2 relates the round trip rate to the rejections, extending Syed et al. (2021, Corollary 2); for the proof, see Appendix B.2.
@@
We now fix $N$ and use Proposition˜2 to study how the round trip rate scales with $K$ in Proposition 3; for the proof, see Appendix B.3. We focus on a special case where ${\mathbb{P}}_{K}^{n-1}$ and ${\mathbb{Q}}_{K}^{n}$ arise as $K$-step discretisations of an underlying stochastic differential equation (SDE) which bridge $\pi^{n-1}$ and $\pi^{n}$ with path measure ${\mathbb{P}}^{n-1}_{\infty}$ and ${\mathbb{Q}}^{n}_{\infty}$ as $K\to\infty$.
@@
Under appropriate conditions on the drifts of the SDE (Appendix B.3), as $K\to\infty$, $\tau({\mathbb{P}}^{0:N-1}_{K},{\mathbb{Q}}^{1:N}_{K})$ converges to $\tau({\mathbb{P}}^{0:N-1}_{\infty},{\mathbb{Q}}^{1:N}_{\infty})$ and $r({\mathbb{P}}_{K}^{n-1},{\mathbb{Q}}_{K}^{n})\leq r({\mathbb{P}}^{n-1}_{\infty},{\mathbb{Q}}^{n}_{\infty})+\mathcal{O}(1/\sqrt{K})$.
The main appeal of APT comes from the fact that for well-designed ${\mathbb{P}}^{0:N-1}_{K},{\mathbb{Q}}^{1:N}_{K}$, that aim to induce approximately the same measure, we have
@@
Thus, increasing $K$ can dramatically reduce the rejection rate and increase the round trip rate in challenging scenarios. In practice, we find that for moderate values of $K$, this benefit outweights the increased computational cost as seen in Section 6.1.
@@
Following the analysis of PT in Syed et al. (2022), we consider how the performance of APT scales as $N\to\infty$. In particular, Theorem˜2 shows that as $N\to\infty$, the round trip rate is controlled by the global barrier $\Lambda_{K}$ for APT and this can be estimated by $\sum_{n=1}^{N}r({\mathbb{P}}^{n-1}_{K},{\mathbb{Q}}^{n}_{K})$ (this can be empirically estimated by approximating $r({\mathbb{P}}^{n-1}_{K},{\mathbb{Q}}^{n}_{K})$ by the empirical rejection rates observed from Algorithm 1); for the proof, see Appendix B.4.2. This provides a direct generalisation of the analysis of the global barrier for PT from Syed et al. (2022).
@@
Suppose ${\mathbb{P}}^{\beta,\beta^{\prime}}_{K}$ and ${\mathbb{Q}}^{\beta,\beta^{\prime}}_{K}$ are sufficiently regular and satisfy Assumptions 2–4 in Appendix B.4. If $\max_{n\leq N}|\beta_{n}-\beta_{n-1}|=O(N^{-1})$ as $N\to\infty$, then $\sum_{n=1}^{N}r({\mathbb{P}}^{n-1}_{K},{\mathbb{Q}}^{n}_{K})$ converges to $\Lambda_{K}$ and $\tau({\mathbb{P}}^{0:N-1}_{K},{\mathbb{Q}}^{1:N}_{K})$ converges to $\bar{\tau}_{K}=(2+2\Lambda_{K})^{-1}$, where $\Lambda_{K}$ equals,
@@
and $\dot{w}^{\beta}_{K}:\mathcal{X}^{K+1}\to\mathbb{R}$ is the partial derivative with respect to $\beta^{\prime}$ of $\log w^{\beta,\beta^{\prime}}_{K}$ at $\beta^{\prime}=\beta$.
Notably, APT observes a sharp deterioration in round trip when $N\ll\Lambda_{K}$, begins to stabilise when $N\approx\Lambda_{K}$ and observes a marginal improvement when $N\gg\Lambda_{K}$. Intuitively, we can consider $\Lambda_{K}$ to be a statistical invariant of APT which quantifies the difficulty of the sampling problem. As discussed in Proposition˜3, for a reasonable choice of accelerators, we expect $\Lambda_{K}$ to decrease with $K$. This suggests APT can substantially improve compared to classical PT ($K=0$) for challenging problems by having $\Lambda_{K}<\Lambda_{0}$.
Schedule tuning.
Another important consequence of Theorem 2 is that this allows us to import the schedule tuning algorithm from (Syed et al., 2022, Section 5) by directly following the same reasoning. This provides a simple and robust method for optimising the annealing schedule $\{\beta_{n}\}_{n=0}^{N}$ which is one of the most important hyper-parameters impacting the performance of APT. We provide further details in Appendix C.
Additional analysis.
Finally, we provide a theoretical equivalence of how APT can be viewed as an instance of PT and explore the trade-off between increasing parallelism and increasing acceleration time in Appendix B.5.
5 Design Space for Accelerated PT
In this section, we explore the design space of APT where $\mathcal{X}=\mathbb{R}^{d}$. We provide several variants of APT algorithms which utilise different neural transports in order to improve the round trip rate of APT.
@@ 5.1 Normalising Flow Accelerated PT
Let $T^{n}:\mathcal{X}\to\mathcal{X}$ be such a diffeomorphism. We define NF-APT as APT with forward and backward accelerators defined by the deterministic transport given by the normalising flow $T^{n}$ and its inverse respectively,
@@
Then the incremental weight equals
$w_{1}^{n}(x_{0},x_{1})=\frac{\tilde{\pi}^{n}(x_{1})}{\tilde{\pi}^{n-1}(x_{0})}|\det J_{T^{n}}(x_{0})|,\quad x_{1}=T^{n}(x_{0}).$
We can parametrise the normalising flow $T^{n}$ by a neural network which is trained to minimise $\mathcal{L}(T)=\sum_{n=1}^{N}\mathrm{SKL}({\mathbb{P}}^{n-1}_{1},{\mathbb{Q}}^{n}_{1})$ following Equation˜4; see Appendix D.1 for further details. We note that NF-APT is partly motivated by prior work (Arbel et al., 2021; Matthews et al., 2022) which utilise normalising flows to map between annealing distributions in the context of SMC samplers (Del Moral et al., 2006). Interestingly, NF-APT is able to mitigate mode dropping issues associated with these works, as we are able to train with the symmetric KL divergence, since NF-APT provides access to samples from both $\pi^{n-1}$ and $\pi^{n}$.
@@ 5.2 Controlled Monte Carlo Diffusions Accelerated PT
Another common transport for sampling is based on the escorted Jarzynski equality (Vaikuntanathan & Jarzynski, 2008). One way to realise this in practice with neural networks is via Controlled Monte Carlo Diffusions (CMCD) (Vargas et al., 2024).
Suppose for $s\in[0,1]$, we have $b^{n}_{s}\in\mathbb{R}^{d}$, $\sigma^{n}_{s}\in\mathbb{R}_{+}$, and $U^{n}_{s}=(1-\phi_{s}^{n})\log\tilde{\pi}^{n-1}+\phi_{s}^{n}\log\tilde{\pi}^{n}$, where $\phi_{s}^{n}\in[0,1]$ is monotonically increasing in $s$ and $\phi^{n}_{0}=0,\phi^{n}_{1}=1$. For a fixed $K$, let $s_{k}=k/K$ and $\Delta s_{k}=1/K$ be the uniform discretisation of the unit interval. We define CMCD-APT as APT with $k$-th forward accelerator,
$P^{n-1}_{k}(x_{k-1},\mathrm{d}x_{k})=\mathcal{N}(x_{k-1}+\Delta s_{k}(\sigma^{n}_{s_{k-1}})^{2}\nabla U^{n}_{s_{k-1}}(x_{k-1})+\Delta s_{k}b^{n}_{s_{k-1}}(x_{k-1}),2\Delta s_{k}(\sigma^{n}_{s_{k-1}})^{2}\mathbf{I}),$
and the backward accelerator,
$Q^{n}_{k-1}(x_{k},\mathrm{d}x_{k-1})=\mathcal{N}(x_{k}-\Delta s_{k}(\sigma_{s_{k}}^{n})^{2}\nabla U^{n}_{s_{k}}(x_{k})+\Delta s_{k}b^{n}_{s_{k}}(x_{k}),2\Delta s_{k}(\sigma_{s_{k}}^{n})^{2}\mathbf{I}).$
The corresponding incremental weights for CMCD-APT equals
$w^{n}_{K}(x_{0:K})=\frac{\tilde{\pi}^{n}(x_{K})\prod_{k=1}^{K}Q^{n}_{k-1}(x_{k},x_{k-1})}{\tilde{\pi}^{n-1}(x_{0})\prod_{k=1}^{K}P^{n-1}_{k}(x_{k-1},x_{k})}$ (5)
where $P^{n-1}_{k}(x_{k-1},x_{k})$ and $Q^{n}_{k-1}(x_{k},x_{k-1})$ are the Gaussian densities of the forward and backwards transition kernels with respect to the Lebesgue measure.
We parametrise and learn the neural transport $b^{n}_{s}$, the time schedule $\phi^{n}_{s}$ and the diffusion coefficient $\sigma^{n}_{s}$ via the objective: $\mathcal{L}(b_{s},\phi_{s},\sigma_{s})=\sum_{n=1}^{N}\mathrm{SKL}({\mathbb{P}}_{K}^{n-1},{\mathbb{Q}}^{n}_{K})$, where we also learn $\phi_{s}^{n},\sigma^{n}_{s}$ following Geffner & Domke (2023) [^3] We note that in standard CMCD, we cannot use the symmetric KL divergence in the objective as standard CMCD only has access to data from one side.. While in theory, once the vectorised field $b_{s}^{n}$ is perfectly learned, any $\sigma^{n}_{s}$ can be used, we found that learning $\sigma^{n}_{s}$ can significantly stabilise the training and enhance performance. We note that alternative divergences and losses can also be employed—see Máté & Fleuret (2023); Richter & Berner (2023); Albergo & Vanden-Eijnden (2024). We discuss the form of the transition kernels, the incremental weights and objective in the limit as $K\to\infty$ in Appendix D.2.
@@ 5.3 Diffusion Accelerated PT
The Variance-Preserving (VP) diffusion model (Ho et al., 2020; Song et al., 2020) is defined by the (forward) SDE: $dY_{s}=-\gamma_{s}Y_{s}\mathrm{d}s+\sqrt{2\gamma_{s}}\mathrm{d}{W}_{s}$ with $s\in[0,1]$, $Y_{0}\sim\pi$, and a rate function $\gamma_{s}:[0,1]\to\mathbb{R}^{+}$. The time-reversal SDE $(X_{s})_{s\in[0,1]}=(Y_{1-s})_{s\in[0,1]}$ has the form $dX_{s}=[\gamma_{1-s}X_{s}+2\gamma_{1-s}\nabla\log\pi_{s}^{\text{VP}}(X_{s})]\mathrm{d}s+\sqrt{2\gamma_{1-s}}\mathrm{d}{W}_{s}$ where $\pi_{s}^{\text{VP}}$ is the density of $Y_{1-s}$ and $X_{0}\sim\pi_{0}^{\text{VP}}$ is close to a standard Gaussian in the common scenario that $\int_{0}^{1}\gamma_{s}\mathrm{d}s\gg 1$. Samples from $\pi$ can then be generated by simulating the time-reversal SDE initialised with samples from $\mathcal{N}(0,\mathbf{I})$; the score $\nabla\log\pi_{s}^{\text{VP}}$ is unknown but can be learned by score matching objectives.
To employ diffusion models for APT, we consider parametrising an energy-based model (Salimans & Ho, 2021) $\pi_{s}^{\theta}$ which satisfies the boundary conditions: $\pi^{\theta}_{0}={\mathcal{N}}(0,\mathbf{I})$ and $\pi^{\theta}_{1}=\pi$, and which we train to approximate $\pi_{s}^{\text{VP}}$. We use $\pi_{s}^{\theta}$ to define our annealing path $\pi^{n}$ by $\pi^{n}=\pi_{s_{n}}^{\theta}$ for some choice of schedule $0=s_{0}<\ldots<s_{N}=1$. Finally, for a fixed $K$, we define Diff-APT as APT with forward and backward accelerator respectively given by some $K$-step integrator of the time-reversal SDE (where we substitute $\pi_{s}^{\theta}$ for $\pi_{s}^{\text{VP}}$) and the forward SDE between the times $s_{n-1}$ and $s_{n}$; for further details, see Appendix D.3.
To provide a concrete example of $P_{k}^{n-1}$ and $Q_{k-1}^{n}$, consider the Euler-Maruyama discretisation with step-size $\delta_{n}=(s_{n}-s_{n-1})/K$ and interpolating times $s_{n}^{k}=s_{n-1}+k\delta_{n}$. The forward and backward accelerators are given by:
$P_{k}^{n-1}(x_{k-1},\mathrm{d}x_{k})$ $=\mathcal{N}(x_{k-1}+\delta_{n}\gamma_{1-s_{n}^{k-1}}[x_{k-1}+2\nabla\log\pi^{\theta}_{s_{n}^{k-1}}(x_{k-1})],2\delta_{n}\gamma_{1-s_{n}^{k-1}}\mathbf{I})$
$Q_{k-1}^{n}(x_{k},\mathrm{d}x_{k-1})$ $=\mathcal{N}(x_{k}-\delta_{n}\gamma_{1-s_{n,k}}x_{k},2\delta_{n}\gamma_{1-s_{n,k}}\mathbf{I}),$
and the resulting formula for the incremental weights $w^{n}_{K}$ is the same as Equation˜5.
Regarding training, as we initially do not have access to samples from $\pi$, we iteratively train, using the score matching objective, on approximate samples from $\pi$ obtained by sampling from Diff-APT. Furthermore, we stress that a parallelised implementation of Diff-APT allows for generating a sample from $\pi$ with a similar effective cost as $K$ discretisation steps solving the time-reversal SDE.